OpenStudy (anonymous):

\[cos^{-1}\left[cos\left(-\frac \pi 3\right)\right]\]

OpenStudy (anonymous):

\[cos\alpha=\frac 12\] \[\alpha =\frac \pi 3\]

zepdrix (zepdrix):

I'm a little confused by the way you wrote that. Perhaps you're taking a different approach than I would :) But it seems to have worked out.

OpenStudy (anonymous):

my ultimate goal is to find alpha correct?

OpenStudy (anonymous):

LOL....I'm sorry

zepdrix (zepdrix):

\[\large \cos^{-1}\left[\cos\left(-\frac \pi 3\right)\right]=\alpha\] Since our outer function is an inverse trig, we're ultimately looking for an angle. Yes I suppose that's true! :) I guess I understand what you did, you just showed the last 2 little steps I guess.

OpenStudy (anonymous):

yep I only showed the last two steps....let's do some sum and difference angle problems :)

OpenStudy (experimentx):